OpenStudy (anonymous):

What are the roots of the function f(x)=(2x-1)(x+5)? A. 1/2 and -5 B. -1 and 5 C. 1 and -5 D. -1/2 and 5

OpenStudy (anonymous):

2x - 1 = 0 ---- 2x = 1 --- x = 1/2 x + 5 = 0 --- x = -5